P=100 TC= 2q+q^2 1) Find q* 2) Find ATC, MC 3) Plot MR, ATC, MC, d...

P=100

TC= 2q+q^2

1) Find q*

2) Find ATC, MC

3) Plot MR, ATC, MC, d and show profit

4) Based on profit what do you expect to happen as a result?
Answer #1

1) Use P = MC. Here MC = 2 + 2q. Then we have 100 = 2 + 2q or q = 98/2 = 49. Hence q* = 49

2) ATC = C/q = 2 + q = $51. MC = 2 + 2*q = 100

3) Graph plot is provided below. Profit is shown as green shaded region

4) We expect new firms entering the industry in the long run due to positive economic profits in short run. This would drive down the price and it will reach the minimum of ATC which is $2.